OpenStudy (anonymous):

solve the system, using substation. Write the solution as an ordered pair (m,n) 2m-3n=6 m=3-2n no solution (-3,-4) (0,-2) (6,2) (3,0)

OpenStudy (anonymous):

I think the solution is the last one (3, 0).

OpenStudy (anonymous):

plug m=3-2n into the first equation \(\Large 2(3-2n)-3n=6\)

OpenStudy (anonymous):

thats what i did and i got n=-4

OpenStudy (anonymous):

Show us your work.

OpenStudy (anonymous):

Did you distribute 2? It would be \(\Large 6-2n-3n=6\)

OpenStudy (anonymous):

It would be 4n not 2n after the 6, sorry

OpenStudy (anonymous):

2(3-2n)-3n=6 6-6n-3n=6 6-3n=6 -3n-6-6 +6 +6 ______________ -3n=12 n=-4 thats my work

OpenStudy (anonymous):

Hmm, In the first step you have 2*-2n=-6n

OpenStudy (anonymous):

When you distribute 2(3-2n) you should get 6-4n because 2*3 = 6 and 2 * -2n = -4n

OpenStudy (anonymous):

flutter you're right.

OpenStudy (anonymous):

Yeah so when you distribute out the paranthesis, you should end up getting an equation that looks like 6-4n-3n = 6. Combine what's called the "like terms" which are -4n and -3n to get -7n. So the equation looks like 6-7n = 6. Then minus 6 from both both sides and you get -7n = 0. Divide by -7 from both sides and n = 0.

OpenStudy (anonymous):

7n-6=6 wouldn't you add?

OpenStudy (anonymous):

wait nevermind i get it

OpenStudy (anonymous):

You flipped the 7n and 6. It's 6-7n not 7n - 6.

OpenStudy (anonymous):

i said i got it